1/3 + (-5/6) + (-1/2) OK I have no idea about these at all. The parenthesis are actually "longer" than a normal parenthesis. I have read all about these and nothing is sinking in for some reason, can someone help me please?
Answer by stanbon(75887)

1/3 + (-5/6) + (-1/2)
--------
The least common denominator is 6:
-------------
= 2/6 -5/6 - 3/6
= (2-5-3)/6
= -6/6
= -1
=============
Cheers,
Stan H.